OpenStudy (anonymous):

Finish writing the equation of the line y=-1/2x+?, given the points (8, -3) and (-2, 2).

OpenStudy (anthonyn2121):

The equation of a slopeis y=mx+b. Because you know a point already, substitute the x coordinate into x and the y coordinate into y. For example, you'd get 2=(1/2)(2)+b.

OpenStudy (anonymous):

okay, how do i find b?

OpenStudy (anthonyn2121):

2=1+b. Subtract 1 and b=1.